Derive mathematical expression of an A.M. wave and indicate the sideband components. Why is the S.S.B. transmission beneficial ?
1
Expert's answer
2019-06-10T04:56:39-0400

Modulation is a technique which modifies waves of high frequency by the wave of low frequency.

Consider the carrier signal:


"V_c(t)=V_c\text{sin}(\omega_ct),"

the modulating signal:


"V_m(t)=V_m\text{sin}(\omega_mt),"

and the signal after modulation look like


"V(t)=[V_c+V_m(t)]\text{sin}(\omega_ct)."

Putting this all together gives


"V(t)=V_c\text{sin}(\omega_ct)+\frac{V_m}{2}\text{cos}(\omega_ct-\omega_mt)-\frac{V_m}{2}\text{cos}(\omega_ct+\omega_mt)."

Now multiply the terms with cosines by "V_c\/V_c":


"V(t)=V_c\text{sin}(\omega_ct)+"

"+\frac{V_m}{2}\frac{V_c}{V_c}\text{cos}(\omega_ct-\omega_mt)-\frac{V_m}{2}\frac{V_c}{V_c}\text{cos}(\omega_ct+\omega_mt),"

and substitute "V_m\/V_c" with "m_a" - the amplitude modulation index.


"V(t)=V_c\text{sin}(\omega_ct)+\frac{V_cm_a}{2}\text{cos}(\omega_ct-\omega_mt)-\frac{V_cm_a}{2}\text{cos}(\omega_ct+\omega_mt)."

The sideband components are


"\frac{\omega_c-\omega_m}{2\pi}=f_c-f_m"

and

"\frac{\omega_c+\omega_m}{2\pi}=f_c+f_m."

The single-sideband transmission reduces the bandwidth transmitted and it allows to double the number of channels within the same frequency band.